The Kelowna division of Windstream RVs builds the Wanderer model. The division had total revenue of $4,785,000 and a profit of $520,000 on the sale of 165 units in the first half of its financial year. Sales declined to 117 units in the second half of the year, resulting in a profit of only $136,000.



Determine the selling price per unit, the total revenue in the second half, the unit variable costs, and the annual fixed costs.
1
Expert's answer
2019-04-16T10:13:23-0400

Price per unit P = $4,785,000/165 = $29,000.

The total revenue in the second half is TR2 = $29,000*117 = $3,393,000.

Total costs in both periods are:

TC1 = TR1 - TP1 = 4,785,000 - 520,000 = $4,265,000,

TC2 = TR2 - TP2 = 3,393,000 - 136,000 = $3,257,000.

TC = FC + AVC×Q -> FC = TC - AVC×Q, so as FC1 = FC2 and AVC1 = AVC2, then:

4,265,000 - 165AVC = 3,257,000 - 117AVC,

AVC = 1,008,000/48 = $21,000 is the unit variable cost.

The annual fixed costs are:

FC1 + FC2 = 2FC2 = 2×(3,257,000 - 117×21,000) = $1,600,000.